OpenStudy (anonymous):

12^{x-2}=20 x=______

OpenStudy (anonymous):

guess my answer didn't help

OpenStudy (anonymous):

no it did, just dont have calculator. but atleast i know how to plug them in lol

OpenStudy (anonymous):

i need the answer in decimal though rounded to the nearest ten-thouandth..

OpenStudy (anonymous):

a calculator? i haven't used a calculator to do anything other than add since 2004 you are on a computer, of course you have a calculator!

OpenStudy (anonymous):

you don't even have to do any work all you have to do is type it in http://www.wolframalpha.com/input/?i=12^ {x-2}%3D20

OpenStudy (anonymous):

U can do that manually

OpenStudy (anonymous):

or type in \[\frac{\ln(20)}{\ln(12)}+2\]

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

can i just give you my username and password and you do it for me?